what is x+5y+4z?
x=7/8, y= 1/9, and z= -1/8

Answer by stanbon(75887)

what is x+5y+4z?
x=7/8, y= 1/9, and z= -1/8
--------------------------------
(7/6) + 5(1/9) + 4(-1/8)
Least common divisor: 72
-----
= (12*7 + 40 -36)/72
-----
= (88/72)
-----
= 11/9
=======================
Cheers,
Stan H.
==============